### v2

# June 4, 2024

Abstract.
Let $P\to X$ be a ${\mathbb{P}}^{1}$-bundle over a variety $X$. The aim of this note is to understand all connected, algebraic groups
$\operatorname{Aut}^{\circ}(P)\subset G\subset\operatorname{Bir}(X\times{\mathbb{P}}^{1}).$
We get a quite complete answer if $\operatorname{Aut}^{\circ}(X)$ is a maximal, connected, algebraic subgroup of $\operatorname{Bir}(X)$, and $P$ is sufficiently unstable.
This gives examples of connected, algebraic subgroups of $\operatorname{Bir}({\mathbb{P}}^{4})$ that are not contained in any maximal one.
Enriques described all connected, algebraic subgroups of the 2-dimensional Cremona group $\operatorname{Bir}({\mathbb{P}}^{2})$ in 1893. In dimension 3, Umemura classified all such subgroups, in a series of papers completed in [Ume88]. They are all contained in a maximal one. Much less is known about algebraic subgroups of $\operatorname{Bir}({\mathbb{P}}^{n})$ for $n\geq 4$.
Fong and Zikas gave examples of connected, algebraic subgroups of $\operatorname{Bir}(C\times{\mathbb{P}}^{n})$ that are not contained in any maximal one, where $C$ is a non-rational curve [Fon24, FZ23]. Fanelli, Floris and Zimmermann constructed similar subgroups in $\operatorname{Bir}({\mathbb{P}}^{n})$ for $n\geq 5$, answering a question of Blanc [BFT23]; see [FFZ23]. The proof in [FFZ23] relies on the existence of a stably rational but nonrational conic bundle 3-fold $X$ [BCTSSD85], a careful investigation of the birational geometry of unstable ${\mathbb{P}}^{1}$-bundles over $X$, and a study of unipotent group actions. For background and previous results, see [BFT22, BFT23] for birational group actions, and [Sar80, Cor95, KM98, KSC04] for minimal models and birational maps.
In this note we follow [Fon24, FZ23, FFZ23], but, instead of using birational geometry, we focus entirely on the action of the unipotent radical of an algebraic group. This leads to examples in $\operatorname{Bir}({\mathbb{P}}^{4})$ as well.
As another consequence, our results apply over algebraically closed fields of arbitrary characteristic.
The main results are Theorems 12 and 20. Here we state a special case that is easy to formulate, yet leads to many interesting subgroups of $\operatorname{Bir}({\mathbb{P}}^{4})$.
Theorem 1.
Let $X$ be a smooth, projective variety such that $\operatorname{Aut}^{\circ}(X)$ is a maximal, connected, algebraic subgroup of $\operatorname{Bir}(X)$. Let $L$ be an $\operatorname{Aut}^{\circ}(X)$-linearized, very ample line bundle on $X$, and
$\operatorname{Aut}^{\circ}\bigl{(}{\mathbb{P}}_{X}({\mathcal{O}}_{X}{\oplus}L)\bigr{)}\subset G\subset\operatorname{Bir}(X\times{\mathbb{P}}^{1})$
a connected, algebraic subgroup. Assume that $(X,L)\not\cong\bigl{(}{\mathbb{P}}^{n},{\mathcal{O}}_{{\mathbb{P}}^{n}}(1)\bigr{)}$.
Then there is an $\operatorname{Aut}^{\circ}(X)$-invariant, effective divisor $D$ on $X$ such that
$G\subset\operatorname{Aut}^{\circ}\bigl{(}{\mathbb{P}}_{X}\bigl{(}{\mathcal{O}}_{X}{\oplus}L(D)\bigr{)}\bigr{)}.$
Note that [Fon24, FZ23, FFZ23] consider cases for which $\operatorname{Aut}^{\circ}(X)=\{1\}$, and use a Noether-Fano type argument to show that $G$ preserves the negative section of ${\mathbb{P}}_{X}({\mathcal{O}}_{X}{\oplus}L)$. It is quite likely that one can prove Theorem 1 using the methods of [Fon24, FZ23, FFZ23], at least in characteristic 0.
For Corollaries 24 we use the notation and assumptions of Theorem 1, and the description of $\operatorname{Aut}^{\circ}\bigl{(}{\mathbb{P}}_{X}({\mathcal{O}}_{X}{\oplus}L)\bigr{)}$ given in Paragraph 18.
Corollary 2.
$\operatorname{Aut}^{\circ}\bigl{(}{\mathbb{P}}_{X}({\mathcal{O}}_{X}{\oplus}L)\bigr{)}$ is a maximal, connected, algebraic subgroup of $\operatorname{Bir}(X\times{\mathbb{P}}^{1})$ iff $H^{0}(X,L)=H^{0}\bigl{(}X,L(D)\bigr{)}$ for every $\operatorname{Aut}^{\circ}(X)$-invariant divisor $D$ on $X$. ∎
Corollary 3.
$\operatorname{Aut}^{\circ}\bigl{(}{\mathbb{P}}_{X}({\mathcal{O}}_{X}{\oplus}L)\bigr{)}$ is a maximal, connected, algebraic subgroup of $\operatorname{Bir}(X\times{\mathbb{P}}^{1})$ for every very ample $L$ iff there are no $\operatorname{Aut}^{\circ}(X)$-invariant divisors on $X$.
Proof. One direction follows from Corollary 2. Conversely, if $D\neq 0$ is an $\operatorname{Aut}^{\circ}(X)$-invariant divisor then there is a very ample $L$ such that $h^{0}(X,L)<h^{0}\bigl{(}X,L(D)\bigr{)}$. Then $\operatorname{Aut}^{\circ}\bigl{(}{\mathbb{P}}_{X}({\mathcal{O}}_{X}{\oplus}L)\bigr{)}\subsetneq\operatorname{Aut}^{\circ}\bigl{(}{\mathbb{P}}_{X}({\mathcal{O}}_{X}{\oplus}L(D))\bigr{)}$
If there are few $\operatorname{Aut}^{\circ}(X)$-orbits, quite likely this can also be proved using the Noether-Fano method [KSC04], but the general case is not clear to me.
Corollary 4.
$\operatorname{Aut}^{\circ}\bigl{(}{\mathbb{P}}_{X}({\mathcal{O}}_{X}{\oplus}L)\bigr{)}$ is not contained in any maximal, connected, algebraic subgroup of $\operatorname{Bir}(X\times{\mathbb{P}}^{1})$, if there is an $\operatorname{Aut}^{\circ}(X)$-invariant divisor $D^{\prime}$ such that $h^{0}\bigl{(}X,{\mathcal{O}}_{X}(D^{\prime})\bigr{)}>1$.
Proof. By Theorem 1, if $\operatorname{Aut}^{\circ}\bigl{(}{\mathbb{P}}_{X}({\mathcal{O}}_{X}{\oplus}L)\bigr{)}\subset G$, then $G\subset\operatorname{Aut}^{\circ}\bigl{(}{\mathbb{P}}_{X}({\mathcal{O}}_{X}{\oplus}L(D))\bigr{)}$ for some $\operatorname{Aut}^{\circ}(X)$-invariant, effective divisor $D$.
Since $h^{0}\bigl{(}X,{\mathcal{O}}_{X}(D^{\prime})\bigr{)}>1$, we see that $h^{0}\bigl{(}X,L(D+D^{\prime})\bigr{)}>h^{0}(X,L(D))$. Thus $\operatorname{Aut}^{\circ}\bigl{(}{\mathbb{P}}_{X}({\mathcal{O}}_{X}{\oplus}L(D))\bigr{)}$ is contained in the strictly bigger group $\operatorname{Aut}^{\circ}\bigl{(}{\mathbb{P}}_{X}({\mathcal{O}}_{X}{\oplus}L(D+D^{\prime}))\bigr{)}$. ∎
Using Umemura’s classification of maximal, connected, algebraic subgroups of $\operatorname{Bir}({\mathbb{C}}{\mathbb{P}}^{3})$, there are 3 cases over ${\mathbb{C}}$ to which Corollary 4 applies. [BFT23] gives an algebraic proof of the classification, which also works in characteristics $p>5$ (though some parts are not yet fully written down).
Example 5.
Let $X$ be a Fano 3-fold with an action $\mathrm{PGL}_{2}\curvearrowright X$ with a dense orbit. The complement is an ample divisor $D$. Thus $\operatorname{Aut}^{\circ}\bigl{(}{\mathbb{P}}_{X}\bigl{(}{\mathcal{O}}_{X}{\oplus}{\mathcal{O}}_{X}(nD)\bigr{)}\bigr{)}$ is not contained in any maximal, connected, algebraic subgroup of $\operatorname{Bir}({\mathbb{P}}^{4})$ for $n\geq 1$.
There are only 2 such examples, both described in [MU83]. One is now called the Mukai-Umemura 3-fold $X\subset{\mathbb{P}}^{21}$. The other is the unique $X\subset{\mathbb{P}}^{6}$ whose general hyperplane section is a degree 5 Del Pezzo surface.
Example 6.
Let $g(u_{0},u_{1})$ be a homogeneous form without multiple roots and
$Q_{g}:=(x_{0}^{2}-x_{1}x_{2}-g(u_{0},u_{1})x_{3}^{2}=0)\subset{\mathbb{P}}^{3}_{x}\times{\mathbb{P}}^{1}_{u}.$
Together with a certain $\mathrm{PGL}_{2}$-action on the ${\mathbb{P}}^{3}_{x}$-factor as in [BFT23, 4.4.4], these are called Umemura quadric fibrations.
We again get that $\operatorname{Aut}^{\circ}\bigl{(}{\mathbb{P}}_{Q_{g}}({\mathcal{O}}_{Q_{g}}{\oplus}L)\bigr{)}$ is not contained in any maximal, connected, algebraic subgroup of $\operatorname{Bir}({\mathbb{P}}^{4})$. The higher dimensional Umemura fibrations studied in [FZ24] give similar examples in $\operatorname{Bir}({\mathbb{P}}^{n})$ for $n>4$.
The subtle example of [FFZ23] can also be treated using Corollary 4.
Example 7.
Let $X$ be a non-rational but rationally connected 3-fold. Then $\{1\}$ is the maximal, connected, algebraic subgroup of $\operatorname{Bir}(X)$. We get that $\operatorname{Aut}^{\circ}\bigl{(}{\mathbb{P}}_{X}({\mathcal{O}}_{X}{\oplus}L)\bigr{)}$ is not contained in any maximal, connected, algebraic subgroup of $\operatorname{Bir}(X\times{\mathbb{P}}^{1})$.
As noted in [FFZ23], there is such an $X$ for which $X\times{\mathbb{P}}^{2}\stackrel{{\scriptstyle bir}}{{\sim}}{\mathbb{P}}^{5}$; see [BCTSSD85]. [FFZ23] uses this to obtain subgroups
$G\cong\operatorname{Aut}^{\circ}\bigl{(}{\mathbb{P}}_{X}({\mathcal{O}}_{X}{\oplus}L)\bigr{)}\times\mathrm{PGL}_{n-3}\subset\operatorname{Bir}({\mathbb{P}}^{n})$
that are not contained in any maximal algebraic subgroup of $\operatorname{Bir}({\mathbb{P}}^{n})$ for $n\geq 5$.
The construction in [FFZ23] uses a non-ample line bundle; these can be treated using Theorem 12.
The next example shows that the assumption $(X,L)\not\cong\bigl{(}{\mathbb{P}}^{n},{\mathcal{O}}_{{\mathbb{P}}^{n}}(1)\bigr{)}$ is necessary in Theorem 1. (Strictly speaking, ${\mathcal{O}}_{{\mathbb{P}}^{n}}(1)$ does not have a $\mathrm{PGL}_{n+1}$-linearization for $n>0$. However, is has an $\mathrm{SL}_{n+1}$-linearization, which is enough for Theorem 1.)
Example 8.
Let $X={\mathbb{P}}^{n}$ and $L={\mathcal{O}}_{{\mathbb{P}}^{n}}(1)$. Then ${\mathbb{P}}_{{\mathbb{P}}^{n}}\bigl{(}{\mathcal{O}}_{{\mathbb{P}}^{n}}{\oplus}{\mathcal{O}}_{{\mathbb{P}}^{n}}(1)\bigr{)}$ is birational to ${\mathbb{P}}^{n+1}$, hence
$\operatorname{Aut}^{\circ}\bigl{(}{\mathbb{P}}_{{\mathbb{P}}^{n}}\bigl{(}{\mathcal{O}}_{{\mathbb{P}}^{n}}{\oplus}{\mathcal{O}}_{{\mathbb{P}}^{n}}(1)\bigr{)}\bigr{)}\subset\operatorname{Aut}^{\circ}({\mathbb{P}}^{n+1})\cong\mathrm{PGL}_{n+2}.$
By contrast, $\operatorname{Aut}^{\circ}\bigl{(}{\mathbb{P}}_{{\mathbb{P}}^{n}}\bigl{(}{\mathcal{O}}_{{\mathbb{P}}^{n}}{\oplus}{\mathcal{O}}_{{\mathbb{P}}^{n}}(r)\bigr{)}\bigr{)}$ is a maximal, connected subgroup of $\operatorname{Bir}({\mathbb{P}}^{n+1})$ for $r\geq 2$ by Corollary 2 (though this is also easy to see directly).
1. Groups acting on ruled varieties
The idea behind the proof could be summarized as follows.
Principle 9.
Let $G\subset\operatorname{Bir}(Y)$ be an algebraic subgroup. If a normal, unipotent subgroup $U\subset G$ acts on $Y$ with small orbits, then every $G\subset\bar{G}\subset\operatorname{Bir}(Y)$ is likely to have a normal, unipotent subgroup $\bar{U}\subset\bar{G}$ that acts on $Y$ with the same orbits.
Theorem 12 illustrates this when the orbits are 1 dimensional.
Definition 10.
Let $\operatorname{Bir}\bigl{(}X\times({\mathbb{P}}^{1},\infty)\to X\bigr{)}\subset\operatorname{Bir}(X\times{\mathbb{P}}^{1})$ denote the subgroup of those birational maps $\phi:X\times{\mathbb{P}}^{1}\dasharrow X\times{\mathbb{P}}^{1}$ that commute with projection to $X$ and map $X\times\{\infty\}$ birationally to $X\times\{\infty\}$. Thus there is a birational map $\phi_{X}:X\dasharrow X$ and a commutative diagram
$\begin{array}[]{ccc}X\times{\mathbb{P}}^{1}&\stackrel{{\scriptstyle\phi}}{{\dasharrow}}&X\times{\mathbb{P}}^{1}\\ \pi\downarrow\hphantom{\pi}&&\hphantom{\pi}\downarrow\pi\\ X&\stackrel{{\scriptstyle\phi_{X}}}{{\dasharrow}}&X,\end{array}$
where $\pi$ is the first coordinate projection.
Note that $\phi\mapsto\phi_{X}$ is a group homomorphism; its kernel is denoted by $\operatorname{Bir}_{X}\bigl{(}X\times({\mathbb{P}}^{1},\infty)\bigr{)}$. Thus we have a group extension
$1\to\operatorname{Bir}_{X}\bigl{(}X\times({\mathbb{P}}^{1},\infty)\bigr{)}\to\operatorname{Bir}\bigl{(}X\times({\mathbb{P}}^{1},\infty)\to X\bigr{)}\to\operatorname{Bir}(X)\to 1$
Definition 11.
Given any connected, algebraic, subgroup $G\subset\operatorname{Bir}\bigl{(}X\times({\mathbb{P}}^{1},\infty)\to X\bigr{)}$, let $G^{v}:=G\cap\operatorname{Bir}_{X}\bigl{(}X\times({\mathbb{P}}^{1},\infty)\bigr{)}$ denote its vertical subgroup, and $G^{h}:=G/G^{v}\subset\operatorname{Bir}(X)$ its horizontal quotient. Thus we have a group extension
$1\to G^{v}\to G\to G^{h}\to 1$
Note that $G^{v}$ is a solvable, normal subgroup of $G$. Let $U^{v}_{G}$ denote its unique maximal, connected, unipotent subgroup. It is a normal subgroup of $G$. Let $U_{G}\subset G$ denote a maximal, connected, unipotent subgroup; it is unique up to conjugation. Note that $U^{v}_{G}\subset U_{G}$, and $U^{h}_{G}:=U_{G}/U^{v}_{G}\subset G^{h}$ is a maximal, connected, unipotent subgroup. (Connectedness is automatic in characteristic 0.)
The following is the main result about subgroups of $\operatorname{Bir}(X\times{\mathbb{P}}^{1})$.
Theorem 12.
Let $X$ be a normal, projective variety and $G\subset\operatorname{Bir}\bigl{(}X\times({\mathbb{P}}^{1},\infty)\to X\bigr{)}$ a connected, algebraic subgroup. Assume that
  1. (12.1)
    $G^{h}\subset\operatorname{Bir}(X)$ is a maximal, connected, algebraic subgroup,
  2. (12.2)
    $\dim U^{v}_{G}\geq 2$, and
  3. (12.3)
    for every connected, algebraic subgroup $\{1\}\neq W\subset U^{h}_{G}$, the fixed point set of the $W$-action by conjugation on $U^{v}_{G}$ has codimension $\geq 2$ in $U^{v}_{G}$.
Then every connected, algebraic subgroup
$G\subset\bar{G}\subset\operatorname{Bir}(X\times{\mathbb{P}}^{1})$
is contained in $\operatorname{Bir}\bigl{(}X\times({\mathbb{P}}^{1},\infty)\to X\bigr{)}$. Furthermore
  1. (12.4)
    $\bar{G}^{h}=G^{h}$, so $U^{h}_{\bar{G}}=U^{h}_{G}$,
  2. (12.5)
    $G^{v}\subset\bar{G}^{v}$, and
  3. (12.6)
    $U^{v}_{G}\subset U^{v}_{\bar{G}}$.
In particular, $\bar{G}$ also satisfies (12.1–3).
Note. (12.3) implies (12.2) whenever $\{1\}\neq U^{h}_{G}$.
Proof. We may assume that $G\neq\bar{G}$.
First we show that there is an intermediate group $G\subsetneq G_{1}\subset\bar{G},$ such that $G_{1}\subset\operatorname{Bir}\bigl{(}X\times({\mathbb{P}}^{1},\infty)\to X\bigr{)}$.
To see this, we apply the purely group theoretic Lemma 16 with $R:=U^{v}_{G}$. Assumption (12.3) says that the alternative (16.2) does not happen.
Thus there is a subgroup $\{1\}\neq R^{\prime}\subset U^{v}_{G}$ whose (connected) normalizer $N_{\bar{G}}(R^{\prime})^{\circ}$ is not contained in $G$.
A key point is that generically $R^{\prime}$ and $U^{v}_{G}$ have the same orbits. In particular, the closures of the 1-dimensional $R^{\prime}$-orbits are still the fibers of the projection $\pi:X\times{\mathbb{P}}^{1}\to X$. Thus $N_{\bar{G}}(R^{\prime})^{\circ}$ birationally preserves $\pi:X\times{\mathbb{P}}^{1}\to X$. The section $X\times\{\infty\}$ is the fixed point set of the $R^{\prime}$-action, so this is also preserved. Thus
$N_{\bar{G}}(R^{\prime})^{\circ}\subset\operatorname{Bir}\bigl{(}X\times({\mathbb{P}}^{1},\infty)\to X\bigr{)}.$
We can then take $G_{1}:=\langle G,N_{\bar{G}}(R^{\prime})^{\circ}\rangle\subset\bar{G}$.
Now $G_{1}$ clearly satisfies $G^{v}\subset G_{1}^{v}$, $U^{v}_{G}\subset U^{v}_{G_{1}}$, and $G^{h}\subset G_{1}^{h}\subset\operatorname{Bir}(X)$. Since $G^{h}$ is a maximal, connected, algebraic subgroup, we must have $G^{h}=G_{1}^{h}$. Therefore $G_{1}$ also satisfies (12.1–3), and induction on $\dim\bar{G}-\dim G$ completes the proof. ∎
Remark 13.
If we do not assume that $G^{h}$ is maximal, the above proof still gives a $G\subsetneq G_{1}\subset\bar{G}$ that is a subgroup of $\operatorname{Bir}\bigl{(}X\times({\mathbb{P}}^{1},\infty)\to X\bigr{)}$. However, if $U_{G}\neq U_{G_{1}}$ then (12.3) may not hold for $G_{1}$, and the induction may break down.
2. Lemmas on algebraic groups
Notation 14.
For an algebraic group $G$, we let $G^{\circ}$ denote the maximal, connected subgroup. $B_{G}\subset G$ denotes a Borel subgroup, $U_{G}\subset B_{G}$ a maximal, connected, unipotent subgroup, and $R_{G}\subset U_{G}$ the unipotent radical of $G$. $Z(\ )$ denotes the center, and $N_{G}(\ )$ the normalizer in $G$.
Lemma 15.
Let $G\subsetneq G^{\prime}$ be connected, algebraic groups. Then there is a subgroup $V\subset U_{G}$ of codimension $\leq 1$ such that $N_{G^{\prime}}\bigl{(}Z(V)\bigr{)}^{\circ}\not\subset G$.
Proof. Let $U_{G}\subset U_{G^{\prime}}$ be maximal unipotent subgroups.
If $U_{G}\neq U_{G^{\prime}}$ then let $U_{G}\subset N\subset U_{G^{\prime}}$ be the normalizer of $U_{G}$ in $U_{G^{\prime}}$. Note that $U_{G}\neq N$ since $U_{G^{\prime}}$ is nilpotent, and $N\not\subset G$ since $U_{G}$ is a maximal unipotent subgroup of $G$. Then $N\subset N_{G^{\prime}}(Z(U_{G}))^{\circ}$, and we can take $V:=U_{G}$.
Otherwise $U_{G}=U_{G^{\prime}}$. For every simple root of $G^{\prime}$ there is a minimal parabolic subgroup $B_{G^{\prime}}\subsetneq P_{i}\subset G^{\prime}$; let $V_{i}\subset U_{G}$ be its unipotent radical. Then $P_{i}/V_{i}$ is either $\mathrm{SL}_{2}$ or $\mathrm{PSL}_{2}$. Therefore $V_{i}\subset U_{G^{\prime}}=U_{G}$ has codimension 1 and $P_{i}\subset N_{G^{\prime}}(V_{i})$.
The $P_{i}$ generate $G^{\prime}$, so $P_{i}\not\subset G$ for some $i_{0}$. Then $V:=V_{i_{0}}$ works. ∎
Lemma 16.
Let $G\subsetneq G^{\prime}$ be connected, algebraic groups, $U_{G}\subset G$ a maximal, connected, unipotent subgroup and $\{1\}\neq R\subset U_{G}$ a connected, normal subgroup. Then one of the following holds.
  1. (16.1)
    There is a connected subgroup $\{1\}\neq R^{\prime}\subset R$ such that $N_{G^{\prime}}(R^{\prime})^{\circ}\not\subset G$.
  2. (16.2)
    There is a connected subgroup $\{1\}\neq W\subset U_{G}/R$ such that the $W$-action by conjugation on $R$ is trivial on a subgroup $V_{R}\subset R$ of codimension $\leq 1$ in $R$.
Proof. By (15) there is a subgroup $V\subset U_{G}$ of codimension $\leq 1$ such that $N_{G^{\prime}}\bigl{(}Z(V)\bigr{)}^{\circ}\not\subset G$. We are done if $Z(V)\subset R$. Otherwise
$W:=\langle Z(V),R\rangle/R\cong Z(V)/\bigl{(}Z(V)\cap R\bigr{)}$
acts trivially on $V_{R}:=V\cap R$, which has codimension $\leq 1$. ∎
3. Regularizing group actions
First recall some results on automorphisms of vector bundles; see for example [Mar71] or [FFZ23, Sec.3].
17Automorphisms of vector bundles.
Let $C$ be a smooth, projective curve and $E$ a rank 2 vector bundle over $C$. Up to twisting with a line bundle, we have one of the following cases.
  1. (17.1)
    $E$ is stable, and then $\operatorname{Aut}_{C}(E)\cong{\mathbb{G}}_{m}$.
  2. (17.2)
    $E\cong{\mathcal{O}}_{C}{\oplus}{\mathcal{O}}_{C}$, and then $\operatorname{Aut}_{C}(E)\cong\mathrm{GL}_{2}$.
  3. (17.3)
    There is an extension $0\to L\to E\to{\mathcal{O}}_{C}\to 0$, where $\deg L\geq 0$ and the extension is non-split if $L\cong{\mathcal{O}}_{C}$. Then the unipotent radical of $\operatorname{Aut}(E)$ is naturally isomorphic to $H^{0}(C,L)$, and
    1. (17.3.a)
      $\operatorname{Aut}_{C}(E)/H^{0}(C,L)\cong{\mathbb{G}}_{m}$ if the extension is non-split, and
    2. (17.3.b)
      $\operatorname{Aut}_{C}(E)/H^{0}(C,L)\cong{\mathbb{G}}_{m}^{2}$ if the extension is split.
Let now $\bar{X}$ be a normal, projective variety and $X\subset\bar{X}$ an open subset such that $\bar{X}\setminus X$ has codimension $\geq 2$. Let $E$ be a rank 2 vector bundle over $X$ such that $\operatorname{Aut}(E)$ contains a unipotent subgroup of dimension $\geq 2$. Working on general curve sections and using (17.1–3) we get that, after twisting with a line bundle and possibly shrinking $X$, there is a unique extension
(17.4)17.4( .4 ) $0\to L\to E\to{\mathcal{O}}_{X}\to 0$
such that $h^{0}(X,L)\geq 2$. Furthermore
  1. (17.5)
    the extension (17.4) is $\operatorname{Aut}_{X}(E)$ invariant,
  2. (17.6)
    the unipotent radical of $\operatorname{Aut}_{X}(E)$ is naturally isomorphic to $H^{0}(X,L)$,
  3. (17.7)
    $\operatorname{Aut}_{X}(E)/H^{0}(X,L)\cong{\mathbb{G}}_{m}$ if (17.4) is non-split, and
  4. (17.8)
    $\operatorname{Aut}_{X}(E)/H^{0}(X,L)\cong{\mathbb{G}}_{m}^{2}$ if (17.4) is split.
18Automorphisms of unstable ${\mathbb{P}}^{1}$-bundles.
Continuing with the notation of (17), let $\pi:P:={\mathbb{P}}_{X}(E)\to X$ be the corresponding ${\mathbb{P}}^{1}$-bundle. Then $\operatorname{Aut}_{X}(P)$ is the quotient if $\operatorname{Aut}_{X}(E)$ by the scalars. Thus we get that, if $\operatorname{Aut}_{X}(P)$ contains a unipotent subgroup of dimension $\geq 2$, then
  1. (18.1)
    there is a unique $\operatorname{Aut}_{X}(P)$ invariant section $D\subset P$,
  2. (18.2)
    the unipotent radical of $\operatorname{Aut}_{X}(P)$ is naturally isomorphic to $H^{0}(X,L)$,
  3. (18.3)
    $\operatorname{Aut}_{X}(P)=H^{0}(X,L)$ if (17.4) is non-split,
  4. (18.4)
    $\operatorname{Aut}_{X}(P)/H^{0}(X,L)\cong{\mathbb{G}}_{m}$ if (17.4) is split, and
  5. (18.5)
    the splittings of $E\to{\mathcal{O}}_{X}$ correspond to the splittings of $\operatorname{Aut}_{X}(P)\to{\mathbb{G}}_{m}$.
Note that the natural way to get a vector bundle on $X$ is to take $\pi_{*}{\mathcal{O}}_{P}(D)$. This sits in the exact sequence
(18.6)18.6( .6 ) $0\to{\mathcal{O}}_{X}=\pi_{*}{\mathcal{O}}_{P}\to\pi_{*}{\mathcal{O}}_{P}(D)\to{\mathcal{O}}_{P}(D)|_{D}\to 0$
Thus $\pi_{*}{\mathcal{O}}_{P}(D)\cong E\otimes L^{-1}$.
Next we show a converse to these observations.
Proposition 19.
Let $\bar{X}$ be a normal, projective variety with smooth locus $X\subset\bar{X}$, and $G\subset\operatorname{Bir}\bigl{(}X\times({\mathbb{P}}^{1},\infty)\to X\bigr{)}$ an algebraic subgroup. Assume that $G^{h}\subset\operatorname{Aut}(\bar{X})=\operatorname{Aut}(X)$. Then there is a $G^{h}$-invariant open subset $X^{\circ}\subset X$, and a ${\mathbb{P}}^{1}$-bundle $\pi^{\circ}:P^{\circ}\to X^{\circ}$ with a rational section $D^{\circ}\subset P^{\circ}$, such that
$G\subset\operatorname{Aut}\bigl{(}(P^{\circ},D^{\circ})\to X^{\circ}\bigr{)}.$
Moreover $X\setminus X^{\circ}$ can be chosen to have codimension $\geq 3$ in $X$.
Theorem 20.
Using the notation and assumptions of (19), assume in addition that there is a ${\mathbb{G}}_{m}\cong T\subset G^{v}/U^{v}_{G}$ and $\dim U^{v}_{G}\geq 1$. Then there is a unique ${\mathbb{P}}^{1}$-bundle $\pi^{m}:P^{m}\to X$ with a $G$-invariant section $D^{m}\subset P^{m}$, such that
  1. (20.1)
    the set of fibers with trivial $U^{v}_{G}$-action has codimension $\geq 2$, and
  2. (20.2)
    for every other $\pi^{\prime}:(P^{\prime},D^{\prime})\to X^{\prime}$ satisfying $G\subset\operatorname{Aut}\bigl{(}(P^{\prime},D^{\prime})\to X^{\prime}\bigr{)}$, there is a natural $G$-equivariant
    $\pi^{\prime}_{*}{\mathcal{O}}_{P^{\prime}}(D^{\prime})\lhook\joinrel\to\pi^{m}_{*}{\mathcal{O}}_{P^{m}}(D^{m}).$
Proof of (19) and (20). By [Wei55] the $G$-action is regularized on some $Y_{1}\stackrel{{\scriptstyle bir}}{{\sim}}\bar{X}\times{\mathbb{P}}^{1}$. Let $Y_{2}\subset\bar{X}\times Y_{1}$ be the closure of the graph of the rational map $Y_{1}\dasharrow\bar{X}\times{\mathbb{P}}^{1}\to\bar{X}$. Then $G$ acts regularly on $Y_{2}$. After normalization we get a $G$-equivariant $Y_{3}\to\bar{X}$, where $Y_{3}$ is normal. It is a ${\mathbb{P}}^{1}$-bundle over a $G^{h}$-invariant, dense open subset $X_{3}\subset X$.
Let $x_{i}\in X\setminus X_{3}$ be the generic points that have codimension 1 in $X$. Working over the localization ${\mathcal{O}}_{x_{i},X}$, (23) gives a $G$-invariant ${\mathbb{P}}^{1}$-bundle extending $Y_{3}\to X_{3}$ over some open neighborhood of $x_{i}$. Gluing these to $Y_{3}\to X_{3}$ we get a $G$-equivariant ${\mathbb{P}}^{1}$-bundle $\pi_{4}:Y_{4}\to X_{4}$ with a rational section $D_{4}\subset Y_{4}$, where $X\setminus X_{4}$ has codimension $\geq 2$ in $X$.
Let $F$ denote the push-forward of $(\pi_{4})_{*}{\mathcal{O}}_{Y_{4}}(D_{4})$ to $X$. Let $X^{\circ}\subset X$ be the largest open subset where $F$ is locally free. Then $X\setminus X^{\circ}$ has codimension $\geq 3$ in $X$.
Set $P^{\circ}:={\mathbb{P}}_{X^{\circ}}\bigl{(}F|_{X^{\circ}}\bigr{)}$. The $G$-action extends to a regular $G$-action on $P^{\circ}$ by by the Matsusaka-Mumford lemma; see [Kol23, 11.39] for the version that we need. This completes the proof of (19).
In (20) the $T$-action on $(\pi_{4})_{*}{\mathcal{O}}_{Y_{4}}(D_{4})$ produces a splitting $(\pi_{4})_{*}{\mathcal{O}}_{Y_{4}}(D_{4})\cong{\mathcal{O}}_{X_{4}}\oplus L^{-1}_{4},$ and $U^{v}_{G}$ is naturally identified with a subspace $W_{G}\subset H^{0}(X,L_{4})$. Let $L_{5}\subset L_{4}$ be the invertible subsheaf generated by the sections in $W_{G}$ at codimension 1 points. Then $L_{5}$ extends to a line bundle $L$ on $X$, and $G$ also acts on $E:={\mathcal{O}}_{X}\oplus L^{-1}$. Its projectivization gives the reguired ${\mathbb{P}}^{1}$-bundle for (20).
Uniqueness can be checked on curves. We can thus work with a rank 2 free module $E_{R}:=Re_{1}\oplus Re_{2}$ over a DVR $R$ with maximal ideal $(t)$ and quotient field $K$, where the $T$ action is $(e_{1},e_{2})\mapsto(\lambda e_{1},\lambda^{-1}e_{2})$, and the elements of the unipotent subgroup act as
$(r_{1}e_{1},r_{2}e_{2})\mapsto\bigl{(}r_{1}e_{1},(r_{2}+ur_{1})e_{2}\bigr{)}\quad\mbox{where}\quad u\in R.$
The unipotent action is nontrivial on the central fiber iff not all the $u$ are divisible by $t$.
Every other $T$-equivariant extension of $E_{K}$ is of the form $E^{*}_{R}:=t^{c_{1}}Re_{1}\oplus t^{c_{2}}Re_{2}$. In the basis $f_{i}:=t^{c_{i}}e_{i}$ the unipotent action becomes
$\bigl{(}r_{1}f_{1},r_{2}f_{2}\bigr{)}\mapsto\bigl{(}r_{1}e_{1},(r_{2}+ut^{c_{1}-c_{2}}r_{1})f_{2}\bigr{)}.$
This is a regular action iff $c_{1}\geq c_{2}$ and it is nontrivial on the central fiber iff $c_{1}=c_{2}$. In this case the projectivizations of $E_{R}$ and of $E^{*}_{R}$ are naturally isomorphic. ∎
Remark 21.
If $\bar{X}$ is factorial, then $P^{m}$ in (20) extends to a ${\mathbb{P}}^{1}$-bundle. If $\bar{X}$ is ${\mathbb{Q}}$-factorial, then $L$ extends to a ${\mathbb{Q}}$-line bundle $\bar{L}$, and $P^{m}$ extends to an orbifold ${\mathbb{P}}^{1}$-bundle
$\operatorname{Proj}_{\bar{X}}\Bigl{(}\oplus_{r\geq 0}\bigl{(}\oplus_{i=0}^{r}\bar{L}^{[i]}\bigr{)}\Bigr{)}.$
Example 22.
The codimension $3$ subset $X\setminus X^{\circ}$ is usually unavoidable in (19), as shown by the following example taken from [AW98]. Consider the second coordinate projection
$\pi:Y:=(\textstyle{\sum}_{i}x_{i}y_{i}=0)\subset{\mathbb{P}}^{2}_{y}\times{\mathbb{A}}^{3}_{x}\to{\mathbb{A}}^{3}_{x}.$
The fiber over $(0,0,0)$ is ${\mathbb{P}}^{2}$, all other fibers are ${\mathbb{P}}^{1}$. We can chose $D:=(y_{0}=0)$ as a rational section. Note that $Y$ is smooth and the projection $\pi$ is $\mathrm{GL}_{3}$-equivariant for the action $\bigl{(}{\mathbf{x}},{\mathbf{y}}\bigr{)}\mapsto\bigl{(}A^{t}{\mathbf{x}},A^{-1}{\mathbf{y}}\bigr{)}.$ (This shows that [FFZ23, Lem.2.2] is not quite true.)
Lemma 23.
Let $k$ be a field, $(t\in T)$ the spectrum of an excellent DVR over $k$, and $\pi:P\to T$ a projective morphism with general fiber ${\mathbb{P}}^{1}$, and $D\subset P$ a section. Let $B$ be the group of $k$-automorphisms of $\pi:(P,D)\to T$.
Then $\pi:(P,D)\to T$ is $B$-equivariantly birational to a trivial bundle $\pi^{\prime}:({\mathbb{P}}^{1}\times T,\{\infty\}\times T)\to T$.
Proof. We can resolve the singularities by normalizing and blowing up the singular set, hence the $B$ action the lifts. Thus we may assume that $P$ is regular. Note that $(-1)$-curves in the central fiber are disjoint from each other, we can thus contract all $(-1)$-curves disjoint from $D$ in a $B$-equivariant way. After such contractions we get a minimal conic bundle with a section, hence a ${\mathbb{P}}^{1}$-bundle. (The residue field of $T$ may be imperfect, see [Tan18] for a proof in these cases.) ∎
Example 24.
The non-uniqueness without the ${\mathbb{G}}_{m}$-action is illustrated by the following, at the vector bundle level.
Let $R$ be a DVR with maximal ideal $(t)$. Consider the rank 2 free module $E:=Re_{1}{\oplus}Re_{2}$ with an endomorphism given by $\phi(e_{1})=e_{2},\phi(e_{2})=0$. The $\phi$-action is nontrivial on $E/tE$. Another extension of $Re_{1}$ by $Re_{2}$ is given by $E^{\prime}:=R(e_{1}+t^{-1}e_{2}){\oplus}Re_{2}$. The $\phi$-action is again nontrivial on $E^{\prime}/tE^{\prime}$.
4. Proof of Theorem 1
25Proof of Theorem 1.
First we aim to apply (12). Note that (12.1) is assumed in Theorem 1. There is nothing to prove if $X$ is a point, otherwise (12.2) also holds. We check in (26) that (12.3) is satisfied. Thus we get that
$G\subset\operatorname{Bir}\bigl{(}X\times({\mathbb{P}}^{1},\infty)\to X\bigr{)}.$
Now we can apply (20) to get that
$G\subset\operatorname{Aut}^{\circ}\bigl{(}{\mathbb{P}}_{X}\bigl{(}{\mathcal{O}}_{X}{\oplus}L_{G}\bigr{)}\bigr{)}$
for some line bundle $L_{G}$.
Set $G_{0}:=\operatorname{Aut}^{\circ}\bigl{(}{\mathbb{P}}_{X}\bigl{(}{\mathcal{O}}_{X}{\oplus}L\bigr{)}\bigr{)}$. Since $L$ is generated by its global sections, ${\mathbb{P}}_{X}({\mathcal{O}}_{X}{\oplus}L)\to X$ is isomorphic to the regularization of $G_{0}$ that is denoted by $P^{m}$ in (20.2). The inclusion in (20.2) induces a $G$-equivariant $L_{G}^{-1}\lhook\joinrel\to L^{-1}$, thus $L_{G}=L(D)$ for some effective $G^{h}$-invariant divisor $D$. ∎
The following accounts for the exceptional case $(X,L)\not\cong\bigl{(}{\mathbb{P}}^{n},{\mathcal{O}}_{{\mathbb{P}}^{n}}(1)\bigr{)}$ in Theorem 1.
Lemma 26.
Let $X$ be a normal, projective variety with a nontrivial ${\mathbb{G}}_{a}$-action. Let $L$ be a very ample, ${\mathbb{G}}_{a}$-linearized line bundle. Assue that the induced ${\mathbb{G}}_{a}$-action is trivial on a codimension 1 subspace $W\subset H^{0}(X,L)$.
Then $X$ is a cone, and the ${\mathbb{G}}_{a}$-action is along its generators.
If $X$ is smooth of dimension $n$, then $(X,L)\cong\bigl{(}{\mathbb{P}}^{n},{\mathcal{O}}_{{\mathbb{P}}^{n}}(1)\bigr{)}$.
Proof. Let $C\subset X$ be the closure of a ${\mathbb{G}}_{a}$-orbit and $c\in C$ the fixed point. $L$ gives an imbedding $X\lhook\joinrel\to{\mathbb{P}}^{N}$; let $\langle C\rangle\subset{\mathbb{P}}^{N}$ be the linear span of $C$. Let $\pi:\bar{C}\to C$ be the normalization. Then
$\dim\operatorname{im}\bigl{[}H^{0}(X,L)\to H^{0}(\bar{C},\pi^{*}L)\bigr{]}=1+\dim\langle C\rangle.$
The ${\mathbb{G}}_{a}$-action on $H^{0}\bigl{(}{\mathbb{P}}^{1},{\mathcal{O}}_{{\mathbb{P}}^{1}}(d)\bigr{)}$ is trivial only on a 1-dimensional subspace. Thus the image of $W$ in $H^{0}(\bar{C},\pi^{*}L)$ has dimension 1 and codimension 1. Therefore $\dim\langle C\rangle=1$, thus $C$ is a line, and the image of $W$ consists of the sections that vanish at $c$.
Given 2 orbits $C_{i}\ni c_{i}$, if $c_{1}\neq c_{2}$ then vanishing at both gives a subset of codimension $\geq 2$. Thus all ${\mathbb{G}}_{a}$-orbit closures are lines, and all these lines pass trough the same point. So $X$ is a cone.
The only smooth cone is a linear subspace. ∎
Acknowledgments.
The main impetus came from the lecture On Algebraic Subgroups of the Cremona Group of Floris at the Conference on Higher Dimensional Geometry (May, 2024) at the Simons Center for Mathematics & Physical Sciences.
In the lecture Floris explained the main steps very clearly, but omitted the technical details. So, when I was trying to think of a proof after the talk, I was free to start on a somewhat different path.
I thank E. Floris and P. Fong for helpful comments. Partial financial support was provided by the NSF under grant number DMS-1901855.